Calculate the surface area of this triangular prism.
A right triangular prism is shown. The right triangular face has base of twelve meters, height of nine meters, and hypotenuse of fifteen meters. The height of the prism is ten meters.

Answer by Alan3354(69443)

Calculate the surface area of this triangular prism.
A right triangular prism is shown. The right triangular face has base of twelve meters, height of nine meters, and hypotenuse of fifteen meters. The height of the prism is ten meters.
-------------------
Nothing is shown.
The triangle is a right triangle, 9, 12 & 15.
The triangle's area = b*h/2 = 9*12/2 = 54 sq meters each end.
------------
The 3 rectangles are 10 by 9, 12 & 15
Area of the 3 = 10*(9 + 12 + 15) = 360 sq meters
-----
Total area = 360 + 2*54
= 468 sq meters